If a pendulum is suspended from a car moving along a plane of angle δ with horizon and car moves with acceleration a, then what is the time period for that pendulum?
1
Expert's answer
2019-03-14T17:45:54-0400
"T=2\pi \sqrt{\dfrac{L}{w}}"

"\vec{w}=\vec{a}+\vec{g}, \,\, w=\sqrt{a^2\cos^2 \delta +g^2}"

"T=2\pi \sqrt{\dfrac{L}{\sqrt{a^2\cos^2 \delta +g^2}}}"

"L" is lenth of pendulum, "w" is resulting acceleration, "g=9.8", "T" is time period.